OpenStudy (anonymous):

Express the fraction in simplest form. 9/12 A. 2/3 B. 3/4 C. 12/16 D. 10/13

OpenStudy (mathmath333):

what is the common multiple between 9 and 12

OpenStudy (anonymous):

36?

OpenStudy (mathmath333):

no

OpenStudy (anonymous):

im not very good at this

OpenStudy (mathmath333):

hint \(3\times 3= 9\\ 3\times 4 =12\)

OpenStudy (anonymous):

:/im not very good at this hint: (mathematics):P

OpenStudy (mathmath333):

lol

OpenStudy (anonymous):

its wierd tho like in the day i dont see all the classes but at 10:00 i do

OpenStudy (anonymous):

or not classes but subjects

OpenStudy (mathmath333):

do u agree that \(3\times 3= 9\)

OpenStudy (anonymous):

i might say stupid things because its 10 and im tired but yeah i agree

OpenStudy (mathmath333):

3+3+3=? if we add 3 times 3 =?

OpenStudy (anonymous):

i said iknow tha stoof i may be tred but i know 3x3

OpenStudy (anonymous):

lol

OpenStudy (mathmath333):

ok so whats \(\dfrac{3\times 3}{3\times 4}\)

OpenStudy (anonymous):

9 and !@ hint !@ is the shifted versions of 1 and 2

OpenStudy (mathmath333):

what

OpenStudy (anonymous):

its 9 and 12

OpenStudy (anonymous):

i told ya i might say stuoopidd stoof

OpenStudy (mathmath333):

lol

OpenStudy (anonymous):

like i said stupid stuff and i sad might im definatley going to

OpenStudy (mathmath333):

\(\dfrac{9}{12}\\ =\dfrac{3\times 3}{3\times 4}\\ =\dfrac{3}{4}\\ \) the common 3 gets cancelled

OpenStudy (anonymous):

so its B)

OpenStudy (anonymous):

can u a helllllpppppp meh???????????????????????????????????????

OpenStudy (anonymous):

@simsim789

OpenStudy (anonymous):

its B. 9/12 divide each by 3, 9/3=3, and 12/3 is 4, so its 3/4

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

6th its hard because im a bit tired

OpenStudy (anonymous):

oh, i can help

OpenStudy (anonymous):

i'm in 8th

OpenStudy (anonymous):

if you need anymore help in anything just tag me